### v1

# MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models

Abstract
Multimodal Large Language Model (MLLM) relies on the powerful LLM to perform multimodal tasks, showing amazing emergent abilities in recent studies, such as writing poems based on an image. However, it is difficult for these case studies to fully reflect the performance of MLLM, lacking a comprehensive evaluation. In this paper, we fill in this blank, presenting the first MLLM Evaluation benchmark MME [^1] Please email guilinli@stu.xmu.edu.cn to obtain the benchmark dataset. This is the V1 version of MME, which will be updated as MLLMs evolve. An online leaderboard will be released at https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models.. It measures both perception and cognition abilities on a total of 14 subtasks. In order to avoid data leakage that may arise from direct use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed. The concise instruction design allows us to fairly compare MLLMs, instead of struggling in prompt engineering. Besides, with such an instruction, we can also easily carry out quantitative statistics. A total of 10 advanced MLLMs are comprehensively evaluated on our MME, which not only suggests that existing MLLMs still have a large room for improvement, but also reveals the potential directions for the subsequent model optimization.
1 Introduction
The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) OpenAI (2023); Alayrac et al. (2022); Huang et al. (2023); Li et al. (2023c); Driess et al. (2023). It refers to using LLM as a brain to process multimodal information and give reasoning results. Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks Driess et al. (2023); Wu et al. (2023); Shen et al. (2023). The three representative abilities of LLM Zhao et al. (2023a), including In-Context Learning (ICL) Brown et al. (2020), instruction following Touvron et al. (2023), and Chain-of-Thought (CoT) Wei et al. (2022), are also manifested in multimodality. For example, Flamingo Alayrac et al. (2022) turns on multimodal ICL, which can adapt to new tasks by giving a few examples. MiniGPT-4 Zhu et al. (2023) implements GPT-4OpenAI (2023)-like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning. PaLM-E Driess et al. (2023) achieves amazing OCR-free math reasoning via CoT. These emergent abilities of MLLM are exciting and imply that a new dawn has broken in artificial intelligence.
Although these models exhibit surprising conversational capabilities when conducting everyday chats, we still know little about how well they quantitatively perform in various aspects. The existing three common quantitative evaluation manners for MLLMs have their limitations that are difficult to comprehensively evaluate performance. Specifically, the first manner Xu et al. (2022); Dai et al. (2023); Wang et al. (2023) evaluates on existing traditional multimodal datasets, such as image caption Chen et al. (2015) and VQA Goyal et al. (2017); Marino et al. (2019); Lu et al. (2022). However, on the one hand, it may be hard to reflect the emergent abilities of MLLM on these datasets. On the other hand, since the training sets of large models are no longer unified, it is difficult to guarantee that all MLLMs have not used the testing set for training. The second manner Ye et al. (2023) is to collect data for an open-ended evaluation, but either the data is unavailable to public by now Zhao et al. (2023c) or the amount is small (only 50 images) Ye et al. (2023). The third manner focuses on one aspect of MLLMs, such as object hallucination Li et al. (2023d) or adversarial robustness Zhao et al. (2023b), which is powerless to comprehensive evaluation.
In light of these concerns, a new comprehensive evaluation benchmark is urgently needed to match the flourish of MLLMs. We argue that a universal comprehensive evaluation benchmark should have the following four characteristics: (1) It should cover as much as possible, including both perception and cognition abilities. The former refers to recognizing the specific object, such as its existence, count, position, and color. The latter refers to compositing the perception information and the knowledge in LLM to deduce more complex answers. It is obvious that the former is the premise of the latter. (2) Its data or annotations should not come from existing publicly available datasets as much as possible, avoiding the risk of data leakage. (3) Its instructions should be as concise as possible and in line with human cognition. Although instruction design may have a large impact on the output, all models should be tested under the same unified instructions for fair comparison. A good MLLM should be able to generalize to such concise instructions. (4) The responses of MLLM to the instructions should be intuitive and convenient for quantitative analysis. The open-ended answer of MLLM poses significant challenges to the quantization. Existing methods tend to use GPT or manual scoring Li et al. (2023a); Liu et al. (2023); Ye et al. (2023), but there may be problems of inaccuracy and subjectivity.
Figure 1: Diagram of our MME benchmark. It evaluates MLLM from both perception and cognition, including a total of 14 subtasks. Each image corresponds to two questions whose answers are marked yes [Y] and no [N], respectively. The instruction consists of a question followed by "Please answer yes or no". It is worth noting that all instructions are manually designed.
To this end, we collect a comprehensive MLLM Evaluation benchmark, named as MME, which meets the above four characteristics at the same time:
  • MME covers the examination of perception and cognition abilities. Apart from OCR, the perception includes the recognition of coarse-grained and fine-grained objects. The former identifies the existence, count, position, and color of objects. The latter recognizes movie posters, celebrities, scenes, landmarks, and artworks. The cognition includes commonsense reasoning, numerical calculation, text translation, and code reasoning. The total number of subtasks is up to 14, as shown in Fig. 1.
  • All instruction-answer pairs are manually constructed. For the few public datasets involved in our study, we only use images without directly relying on their original annotations. Meanwhile, we make efforts to collect data through real photographs and image generation.
  • The instructions of MME are designed concisely to avoid the impact of prompt engineering on the model output. We argue that a good MLLM should be able to generalize to such simple and frequently used instructions, which are fair to all models. Please see Fig. 1 for the specific instruction of each subtask.
  • Benefitting from our instruction design “please answer yes or no”, we can easily perform quantitative statistics based on the “yes” or “no” output of MLLMs, which is accurate and objective. It should be noted that we have also tried to design instructions with multiple choice questions, but find that it may beyond the capabilities of current MLLMs to follow complex instructions.
We conduct massive experiments to evaluate the zero-shot performance of 10 advanced MLLMs on 14 subtasks. The evaluated MLLMs include BLIP-2 Li et al. (2023c), LLaVA Liu et al. (2023), MiniGPT-4 Zhu et al. (2023), mPLUG-Owl Ye et al. (2023), LLaMA-Adapter-v2 Gao et al. (2023), Otter Li et al. (2023b), Multimodal-GPT Gong et al. (2023), InstructBLIP Dai et al. (2023), VisualGLM-6B git (2023i), and PandaGPT Su et al. (2023). As displayed in Fig. 2 that consists of 2 overall leaderboards (perception and cognition) and 14 individual leaderboards, these MLLMs show clear discrepancies in our MME evaluation benchmark. Fig. 3 also provides a comparison from the other perspective. More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination, as shown in Fig. 4. These findings are instructive for the subsequent model optimization.
In summary, the contributions of this work are as follows: (1) We propose a new benchmark MME to meet the urgent need of MLLM evaluation. As far as we know, MME is the first comprehensive MLLM evaluation benchmark [^2] The concurrent works are LAMM-Benchmark Yin et al. (2023) and LVLM-eHub Xu et al. (2023), which are rather different from our MME. We recommend reading the two works as well to enrich the understanding of MLLM evaluation.. (2) We conduct abundant experiments on a total of 14 subtasks to comprehensively evaluate 10 up-to-date MLLMs. (3) We summarize the exposed problems in experiments, proving guidance for the evolution of MLLM.
2 MME Evaluation Suite
2.1 Instruction Design
In order to facilitate quantitative performance statistics, the orientation of our instruction design is to let the model to answer “yes” or “no”. As a result, the instruction consists of two parts, including a concise question and a description “Please answer yes or no.” For each test image, we manually design two instructions, where the discrepancy lies in the questions. The ground truth answer of the first question is “yes” and that of the second question is “no”, as shown in Fig. 1. When MLLM answers both of these questions correctly, it appears more confident that the MLLM actually comprehends the image and the corresponding knowledge behind it.
2.2 Evaluation Metric
Since the output of the model is limited to two types (“yes” or “no”), it is convenient to measure the metrics of accuracy and accuracy+. The former is calculated based on each question, while the latter is based on each image where both of the two questions need to be answered correctly. The random accuracies of the two metrics are equal to 50% and 25%, respectively. It can be seen that accuracy+ is a stricter measurement but also better reflects the comprehensive understanding degree of the model to the image. In addition, we calculate the score of a subtask based on the sum of accuracy and accuracy+. The perception score is the sum of scores of all perception subtasks. The cognition score is calculated in the same way.
Figure 2: Leaderboards on our MME benchmark. (1) and (2) are the overall leaderboards of perception and cognition respectively, in which the full score of the former is 2000 and that of the latter is 800. (3)-(16) are the leaderboards of the 14 subtasks with the full score of 200. The score is the sum of the accuracy and the accuracy+ in Tables. 1, 2, and 3.
Figure 3: Comparison of 10 advanced MLLMs on 14 subtasks. The full score of each subtask is 200.
2.3 Data Collection
2.3.1 Perception Tasks
We argue that perception is one of the most fundamental capabilities of MLLM, and the lack of perception will easily lead to the object hallucination problem Li et al. (2023d). That is, MLLM will answer questions based on its own fantasies rather than based on the realistic content of the image, as displayed in Fig. 4.
Coarse-Grained Recognition. The contents of coarse-grained recognition include the existence of common objects, and their count, color, and position. The images are sampled from COCO Lin et al. (2014), but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations. Even if MLLMs have seen these COCO images, our manually prepared pairs are not presented in their training sets. This requires MLLM to be able to understand the instructions and infer corresponding answers. In each perception subtask of existence, count, color, and position, we prepare 30 images with 60 instruction-answer pairs.
Fine-Grained Recognition. The fine-grained recognition is more about testing the knowledge resources of MLLM. The subtasks consist of recognizing movie posters, celebrities, scenes, landmarks, and artworks, containing 147, 170, 200, 200, and 200 images respectively. For the celebrities, we plot a red box to a person with a clearly visible face in the image, and the corresponding instruction is “Is the actor inside the red box named [celebrity name]? Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets Huang et al. (2020); Mao et al. (2017, 2019); Zhou et al. (2014); Weyand et al. (2020) and all of the instructions are manually designed.
OCR. Optical Character Recognition (OCR) is also a foundational capability of MLLM, serving for subsequent text-based tasks such as text translation and text understanding. The images are from Liu et al. (2019) and the instruction-answer pairs are manually designed. Considering that MLLM is still in its infancy, we only choose the relatively simple samples in this version of MME. The numbers of image and instruction-answer pairs are 20 and 40, respectively.
2.3.2 Cognition Tasks
We evaluate if any MLLM can carry out further logical reasoning after perceiving the image, which is the most fascinating aspect of MLLM over previous traditional methods. In order to infer the correct answer, MLLM needs to follow the instruction, perceive the contents of the image, and invoke the knowledge reserved in LLM, which is much more challenging than the single perception tasks. Examples of the following subtasks are shown in Fig. 1.
Commonsense Reasoning. Unlike the ScienceQA dataset Lu et al. (2022) that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life. For example, given a photo of a down jacket, asking MLLM whether it is appropriate to wear the cloth when it is cold (or hot). These are basic knowledge that humans can judge instantly without complex step-by-step reasoning. Therefore, we expect MLLM to perform well in a zero-short setting. The images are all manually photographed or generated by diffusion models, and the instruction-answer pairs are all manually designed. There are a total of 70 images and 140 instruction-answer pairs.
Numerical Calculation. It requires MLLM to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in Huang et al. (2023). In this version, we only consider relatively easy arithmetic problems, such as addition, subtraction, multiplication, and division. There are a total of 20 images and 40 instruction-answer pairs. The images are all manually taken, and the instruction-answer pairs are all manually designed.
Text Translation. Considering that the MLLM git (2023i) supports both English and Chinese, we set the text translation task. It requires MLLM to translate the Chinese written in an image to the corresponding English. In this version, we only design basic translation problems, which will be updated according to the development of MLLM in the future. The images of this part are all manually taken, and the instruction-answer pairs are all manually designed. There are a total of 20 images and 40 instruction-answer pairs.
Code Reasoning. It requires MLLM to read the code in the images and automatically complete logical operation inside the code. A similar task that writes website code based on an image has been demonstrated in Zhu et al. (2023). The images are all manually taken, and the instruction-answer pairs are all manually designed. We only set basic code problems in this version. There are in total 20 images and 40 instruction-answer pairs.
3 Experiments
Model Existence Count Position Color OCR ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 86.67 73.33 75.00 60.00 56.67 16.67 81.67 66.67 70.00 40.00 LLaVA 50.00 0.00 50.00 0.00 50.00 0.00 51.67 3.33 50.00 0.00 MiniGPT-4 75.00 60.00 66.67 56.67 56.67 33.33 71.67 53.33 62.50 35.00 mPLUG-Owl 73.33 46.67 50.00 0.00 50.00 0.00 51.67 3.33 55.00 10.00 LLaMA-AdapterV2 76.67 56.67 58.33 6.67 43.33 3.33 55.00 16.67 57.50 15.00 InstructBLIP 95.00 90.00 80.00 63.33 53.33 13.33 83.33 70.00 57.50 15.00 VisualGLM-6B 61.67 23.33 50.00 0.00 48.33 0.00 51.67 3.33 42.50 0.00 Otter 53.33 6.67 50.00 0.00 50.00 0.00 51.67 3.33 50.00 0.00 Multimodal-GPT 46.67 10.00 51.67 6.67 45.00 13.33 55.00 13.33 57.50 25.00 PandaGPT 56.67 13.33 50.00 0.00 50.00 0.00 50.00 0.00 50.00 0.00
Table 1: Evaluation of coarse-grained recognition and OCR.
Model Poster Celebrity Scene Landmark Artwork ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 79.25 62.59 68.53 37.06 81.25 64.00 79.00 59.00 76.50 60.00 LLaVA 50.00 0.00 48.82 0.00 50.00 0.00 50.00 0.00 49.00 0.00 MiniGPT-4 49.32 19.73 58.82 24.71 68.25 45.50 59.75 30.50 56.25 27.00 mPLUG-Owl 77.89 57.14 66.18 34.12 78.00 57.50 86.25 73.00 63.25 33.00 LLaMA-AdapterV2 52.72 10.88 55.00 21.18 68.75 44.50 53.00 9.00 52.50 14.50 InstructBLIP 74.15 49.66 67.06 34.12 84.00 69.00 59.75 20.00 76.75 57.50 VisualGLM-6B 54.42 12.24 50.88 2.35 81.75 64.50 59.75 24.00 55.75 20.00 Otter 45.24 0.00 50.00 0.00 55.00 14.50 52.00 4.50 48.00 5.50 Multimodal-GPT 45.24 17.01 49.12 24.12 50.50 17.50 50.50 23.00 46.00 12.00 PandaGPT 56.80 19.73 46.47 10.59 72.50 45.50 56.25 13.50 50.25 1.00
Table 2: Evaluation of fine-grained recognition.
Model Commonsense Reasoning Numerical Calculation Text Translation Code Reasoning ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 68.57 41.43 40.00 0.00 55.00 10.00 55.00 20.00 LLaVA 49.29 11.43 50.00 0.00 52.50 5.00 50.00 0.00 MiniGPT-4 58.57 34.29 47.50 20.00 42.50 15.00 67.50 45.00 mPLUG-Owl 59.29 24.29 50.0 10.00 60.00 20.00 47.50 10.00 LLaMA-Adapter V2 54.29 14.29 52.50 5.00 52.50 5.00 52.50 10.00 InstructBLIP 75.00 54.29 35.00 5.00 55.00 10.00 47.50 0.00 VisualGLM-6B 45.71 12.86 45.00 0.00 55.00 10.00 50.00 0.00 Otter 48.57 10.00 47.50 10.00 55.00 10.00 50.00 0.00 Multimodal-GPT 45.71 5.71 50.00 20.00 50.00 5.00 45.00 10.00 PandaGPT 56.43 17.14 50.00 0.00 52.50 5.00 47.50 0.00
Table 3: Evaluation of cognition.
In this section, we conduct massive experiments on our MME benchmark to evaluate a total of 10 open-source MLLMs, including BLIP-2, LLaVA, MiniGPT-4, mPLUG-Owl, LLaMA-Adapter-v2, Otter, Multimodal-GPT, InstructBLIP, VisualGLM-6B, and PandaGPT. Except BLIP-2, other models have been fine-tuned on their instruction tuning datasets.
3.1 Models
All experiments are conducted on NVIDIA V100 GPUs, and we use as large models as possible. BLIP-2 Li et al. (2023c) focuses on the basic pre-training, where the image encoder and the LLM are all frozen and a light-weight Q-Former is trained for multimodal representation alignment and vision-to-language generation. Despite not performing the multimodal instruction tuning, BLIP-2 is still able to follow instructions with the LLM’s own capabilities. The version under investigation is “blip2-pretrain-flant5xxl” git (2023a).
LLaVA Liu et al. (2023) is a pioneer in bringing multimodal instruction tuning to MLLM. The tuning data are generated by language-only GPT-4 through carefully designed prompts. The instructions consist of three types, including conversation, detailed description, and complex reasoning. A projection layer and LLM are updated during training. The tested version is “LLaVA-7B-v0” git (2023d).
MiniGPT-4 Zhu et al. (2023) performs multimodal instruction tuning on the pre-trained BLIP-2, where only a linear layer is updated. The instructions are mainly based on the image caption task, such as “Describe this image in detail.” We use “minigpt4-aligned-with-vicuna13b” git (2023e) for testing.
mPLUG-Owl Ye et al. (2023) uses both language-only instruction data and the multimodal instruction data from LLaVA. The visual encoder, the proposed visual abstractor, and LLM (insert a LoRA module Hu et al. (2021)) are all updated during training. We test “mplug-owl-llama-7b” git (2023f).
LLaMA-Adapter V2 Gao et al. (2023) is only trained on language-only instruction data and image-text pairs via a parameter-efficient manner. Additional expert models, such as detection, can be integrated to boost reasoning. The version of “LLaMA-7B” git (2023c) is used in testing.
InstructBLIP Dai et al. (2023) is based on BLIP-2 and reorganizes 26 existing public datasets like image caption and VQA as instruction tuning format. Only the Q-Former is updated during training. “blip2-instruct-flant5xxl” git (2023b) is tested.
VisualGLM-6B git (2023i) is an open-source MLLM but with unclear training details by now. It supports both English and Chinese. We test “VisualGLM-6B”  git (2023i).
Otter Li et al. (2023b) mixes multimodal in-context learning and multimodal instruction tuning. The framework is based OpenFlamingo Awadalla et al. (2023), whose perceiver module and parts of LLM are updated during training. The instruction data are from VQAv2 Antol et al. (2015), GQA Hudson and Manning (2019), LLaVA, and a not public video dataset. The version of “OTTER-9B-LA-InContext” git (2023g) is tested.
Multimodal-GPT Gong et al. (2023) is also based on OpenFlamingo and uses both language-only and multimodal instruction data. The latter includes data from LLaVA, Mini-GPT4, and various image caption and VQA datasets. Only the LoRA module in LLM is updated during training. The tested version is “Multimodal-GPT-9B”.
PandaGPT Su et al. (2023) uses multimodal instruction tuning data from LLaVA and MiniGPT-4, and only a linear projection layer as well as the LoRA module in LLM are trainable. The pretrained ImageBind Girdhar et al. (2023) is adopted as the multimodal encoder, which has an attribute of modality alignment. “pandagpt-7b-max-len-512” git (2023h) is used for testing.
3.2 Results
3.2.1 Perception
There are a total of 10 subtasks for the evaluation of the perception ability, from the perspectives of coarse-grained recognition, fine-grained recognition, and OCR. Figs. 2 (3)-(6) show the score leaderboards of individual coarse-grained recognition subtasks. With respect to the object existence, InstructBLIP gets the highest score 185, with a 95% accuracy and a 90% accuracy+ listed in Table 1. Contrastively, the second place BLIP-2 and the third place MiniGPT-4 lag behind InstructBLIP by 25 and 50 scores, respectively. For the object count, position, and color, InstructBLIP, BLIP-2, and MiniGPT-4 still make the top three, which is reasonable because the object existence is the basis of object attributes. What is interesting is that both InstructBLIP and MiniGPT-4 are instruction finetuned on the pre-trained BLIP-2. Note that in the four coarse-grained subtasks, these MLLMs get the worst results on object position, indicating that the current models are not sensitive enough to the position information.
Figs. 2 (7)-(11) show the score leaderboards of individual fine-grained recognition subtasks. Regarding to poster recognition, BLIP-2, mPLUG-Owl, and InstructBLIP are the top three. It is interesting that mPLUG-Owl underperforms in the coarse-grained recognition, but now it exhibits good. This implies that our division of coarse-grained and fine-grained is reasonable and enables us to examine different aspects of MLLMs. For the celebrity recognition, BLIP-2, InstructBLIP, and mPLUG-Owl still take the top three with similar scores. For the scene recognition, InstructBLIP, VisualGLM-6B, and BLIP-2 ahead of other MLLMs. This is the first time VisualGLM-6B has broken into the top three. Also noteworthy is PandaGPT, which scores more than a hundred for the first time. For the landmark recognition, top three places are taken by mPLUG-Owl, BLIP-2, and MiniGPT-4 respectively, of which mPLUG-Owl gets the top spot. For the artwork recognition, BLIP-2, InstructBLIP, and mPLUG-Owl exceed other counterparts, where the first two scores are similar, leading the third by 38+. With respect to OCR listed in Fig. 2 (12), BLIP-2, MiniGPT-4, and Multimodal-GPT get the top three with scores of 110, 97.5, and 82.5 respectively, in which the first two perform more outstanding.
Figure 4: Common problems revealed in our experiments. [Y]/[N] means the ground truth answer is yes/no. [R] is the generated answer of MLLM.
As can be seen in Fig. 2 (1), in the leaderboard of the whole perception recognition, BLIP-2, InstructBLIP, and MiniGPT-4 come in top three, closely followed by mPLUG-Owl. Their perception scores are 1293.84, 1212.82, 1010.66, and 966.32, respectively. Among them, mPLUG-Owl shows a clear bias, good at fine-grained recognition, but not good at coarse-grained recognition. On the contrary, MiniGPT-4 has worse fine-grained recognition than coarse-grained recognition, while the discrepancy is not that significant as mPLUG-Owl. Comparatively, the overall performances of BLIP-2 and InstructBLIP are more stable.
3.2.2 Cognition
There are four subtasks for the evaluation of the cognition ability, including commonsense reasoning, numerical calculation, text translation, and code reasoning. Figs. 2 (13)-(16) plot the score leaderboards of individual subtasks. In terms of the commonsense reasoning, the “ever-victorious generals” InstructBLIP, BLIP-2, and MiniGPT-4 exceed other MLLMs, especially InstructBLIP, which gets a score of 129.29. With respect to numerical calculation and text translation, although the designed questions are moderate in difficulty, these MLLMs do not perform satisfactorily, with no score beyond 80. This suggests that MLLMs have a lot of room for improvement in these capabilities. For the code reasoning, MiniGPT-4 achieves a high score of 112.5, far ahead of other counterparts. This is reminiscent of the example it showed of writing website codes from an image. For all of the cognition tasks, MiniGPT-4, InstructBLIP, and BLIP-2 win the gold, silver, and bronze medals respectively, as shown in Fig. 2 (2).
4 Analysis
We conclude four common problems that largely affect the performance of MLLMs. The first problem is not following instructions. Although we have adopted a very concise instruction design, there are MLLMs that answer freely rather than following instructions. For example, as shown in the first row of Fig. 4, the instruction has claimed “Please answer yes or no”, but the MLLM only makes a declarative expression. If no “yes” or “no” is appeared at the beginning of the generated languages, the model is judged to make a wrong answer. We argue that a good MLLM (especially after instruction tuning) should be able to follow such a simple instruction, which is also very common in everyday life.
The second problem is a lack of perception. As shown in the second row of Fig. 4, the MLLM misidentifies the number of bananas in the first image, and misreads the characters in the second image, resulting in wrong answers. We notice that the performance of perception is vulnerable to the nuance of instructions, since the two instructions of the same image differ in only one word, but lead to completely different and even contradictory perception results.
The third problem is a lack of reasoning. In the third row of Fig. 4, we can see from the red text that the MLLM already knows that the first image is not an office place, but still gives an incorrect answer of “yes”. Analogously, in the second image, the MLLM has calculated the right arithmetic result, but finally delivers a wrong answer. These phenomena indicate that the logic chain is broken during the reasoning process of MLLMs. Adding CoT prompts, such as “Let’s think step by step” Driess et al. (2023), may yield better results. We look forward to a further in-depth research.
The fourth problem is object hallucination following instructions, which is exemplified in the fourth row of Fig. 4. When the instruction contains descriptions of an object that does not appear in the image, the MLLM will imagine that the object exists and ultimately gives a “yes” answer. Such a case of constantly answering “yes” results in an accuracy about 50% and an accuracy+ about 0, as shown in Tables 1, 2, and 3. This suggests an urgent need to suppress hallucinations, and the community should take into account of the reliability of the generated answers.
5 Conclusion
This paper has presented the first MLLM evaluation benchmark MME that has four distinct characteristics in terms of task type, data source, instruction design, quantitative statistics. We evaluate 10 advanced MLLMs on MME and the experimental results show that there is still a large room to improve. We also summarize the common problem raised in experimental results, providing valuable guidance for the development of MLLM.

### v4

@@
 However, it is difficult for these case studies to fully reflect the performance of MLLM, lacking a comprehensive evaluation.
-In this paper, we fill in this blank, presenting the first MLLM Evaluation benchmark MME [^1] Please email guilinli@stu.xmu.edu.cn to obtain the benchmark dataset.
-This is the V1 version of MME, which will be updated as MLLMs evolve.
-An online leaderboard will be released at https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models..
+In this paper, we fill in this blank, presenting the first comprehensive MLLM Evaluation benchmark MME [^1] MME is collected by Xiamen University for academic research only.
+This is the v1 version and will be updated as MLLMs evolve..
 It measures both perception and cognition abilities on a total of 14 subtasks.
@@
 Besides, with such an instruction, we can also easily carry out quantitative statistics.
-A total of 10 advanced MLLMs are comprehensively evaluated on our MME, which not only suggests that existing MLLMs still have a large room for improvement, but also reveals the potential directions for the subsequent model optimization.
+A total of 30 advanced MLLMs are comprehensively evaluated on our MME, which not only suggests that existing MLLMs still have a large room for improvement, but also reveals the potential directions for the subsequent model optimization.
+The data application manner and online leaderboards are released at https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation.
 1 Introduction
-The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) OpenAI (2023); Alayrac et al.
-(2022); Huang et al.
-(2023); Li et al.
-(2023c); Driess et al.
-(2023).
-It refers to using LLM as a brain to process multimodal information and give reasoning results.
-Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks Driess et al.
-(2023); Wu et al.
-(2023); Shen et al.
-(2023).
-The three representative abilities of LLM Zhao et al.
-(2023a), including In-Context Learning (ICL) Brown et al.
-(2020), instruction following Touvron et al.
-(2023), and Chain-of-Thought (CoT) Wei et al.
-(2022), are also manifested in multimodality.
-For example, Flamingo Alayrac et al.
-(2022) turns on multimodal ICL, which can adapt to new tasks by giving a few examples.
-MiniGPT-4 Zhu et al.
-(2023) implements GPT-4OpenAI (2023)-like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning.
-PaLM-E Driess et al.
-(2023) achieves amazing OCR-free math reasoning via CoT.
-These emergent abilities of MLLM are exciting and imply that a new dawn has broken in artificial intelligence.
+The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) [37, 8, 20, 23, 13, 9].
+It refers to using LLM as a brain to process multimodal information and give reasoning results [49].
+Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks [13, 45, 38, 27, 54, 31].
+The three representative abilities of LLM [55], including instruction following [40], In-Context Learning (ICL) [10], and Chain-of-Thought (CoT) [43] are also manifested in multimodality.
+For example, Flamingo [8] turns on multimodal ICL, which can adapt to new tasks by giving a few examples.
+PaLM-E [13] achieves amazing OCR-free math reasoning via CoT.
+GPT-4V [37] shows even more ability in a variety of complex reasoning tasks [46].
+MiniGPT-4 [59] implements GPT-4[37]-like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning.
+These emergent abilities of MLLMs are exciting and imply that a new dawn has broken in artificial intelligence.
+Figure 1: Diagram of our MME benchmark.
+It evaluates MLLMs from both perception and cognition, including a total of 14 subtasks.
+Each image corresponds to two questions whose answers are marked yes [Y] and no [N], respectively.
+The instruction consists of a question followed by “Please answer yes or no”.
+It is worth noting that all instructions are manually designed.
+Figure 2: Leaderboards on our MME benchmark.
+(1) and (2) are the overall leaderboards of perception and cognition respectively, in which the full score of the former is 2000 and that of the latter is 800.
+(3)-(16) are the leaderboards of the 14 subtasks with the full score of 200.
+The score is the sum of the accuracy and the accuracy+ in Tables.
+1 and 2.
+A total of 30 advanced MLLMs joint the leaderboards.
+For the sake of presentation, we only show 10 models for each list, in which the top three ones are given clear trophy logos.
 Although these models exhibit surprising conversational capabilities when conducting everyday chats, we still know little about how well they quantitatively perform in various aspects.
 The existing three common quantitative evaluation manners for MLLMs have their limitations that are difficult to comprehensively evaluate performance.
-Specifically, the first manner Xu et al.
-(2022); Dai et al.
-(2023); Wang et al.
-(2023) evaluates on existing traditional multimodal datasets, such as image caption Chen et al.
-(2015) and VQA Goyal et al.
-(2017); Marino et al.
-(2019); Lu et al.
-(2022).
-However, on the one hand, it may be hard to reflect the emergent abilities of MLLM on these datasets.
+Specifically, the first manner [47, 12, 42] evaluates on existing traditional multimodal datasets, such as image caption [11] and VQA [16, 36, 32].
+However, on the one hand, it may be hard to reflect the emergent abilities of MLLMs on these datasets.
 On the other hand, since the training sets of large models are no longer unified, it is difficult to guarantee that all MLLMs have not used the testing set for training.
-The second manner Ye et al.
-(2023) is to collect data for an open-ended evaluation, but either the data is unavailable to public by now Zhao et al.
-(2023c) or the amount is small (only 50 images) Ye et al.
-(2023).
-The third manner focuses on one aspect of MLLMs, such as object hallucination Li et al.
-(2023d) or adversarial robustness Zhao et al.
-(2023b), which is powerless to comprehensive evaluation.
+The second manner [48] is to collect data for an open-ended evaluation, but either the data is unavailable to public by now [57] or the amount is small (only 50 images) [48].
+The third manner focuses on one aspect of MLLMs, such as object hallucination [25] or adversarial robustness [56], which is powerless to comprehensive evaluation.
 In light of these concerns, a new comprehensive evaluation benchmark is urgently needed to match the flourish of MLLMs.
@@
 A good MLLM should be able to generalize to such concise instructions.
-(4) The responses of MLLM to the instructions should be intuitive and convenient for quantitative analysis.
-The open-ended answer of MLLM poses significant challenges to the quantization.
-Existing methods tend to use GPT or manual scoring Li et al.
-(2023a); Liu et al.
-(2023); Ye et al.
-(2023), but there may be problems of inaccuracy and subjectivity.
-Figure 1: Diagram of our MME benchmark.
-It evaluates MLLM from both perception and cognition, including a total of 14 subtasks.
-Each image corresponds to two questions whose answers are marked yes [Y] and no [N], respectively.
-The instruction consists of a question followed by "Please answer yes or no".
-It is worth noting that all instructions are manually designed.
+(4) The responses of MLLMs to the instructions should be intuitive and convenient for quantitative analysis.
+The open-ended answer of MLLMs poses significant challenges to the quantization.
+Existing methods tend to use GPT or manual scoring [21, 29, 48], but there may be problems of inaccuracy and subjectivity.
 To this end, we collect a comprehensive MLLM Evaluation benchmark, named as MME, which meets the above four characteristics at the same time:
@@
 It should be noted that we have also tried to design instructions with multiple choice questions, but find that it may beyond the capabilities of current MLLMs to follow complex instructions.
-We conduct massive experiments to evaluate the zero-shot performance of 10 advanced MLLMs on 14 subtasks.
-The evaluated MLLMs include BLIP-2 Li et al.
-(2023c), LLaVA Liu et al.
-(2023), MiniGPT-4 Zhu et al.
-(2023), mPLUG-Owl Ye et al.
-(2023), LLaMA-Adapter-v2 Gao et al.
-(2023), Otter Li et al.
-(2023b), Multimodal-GPT Gong et al.
-(2023), InstructBLIP Dai et al.
-(2023), VisualGLM-6B git (2023i), and PandaGPT Su et al.
-(2023).
+We conduct massive experiments to evaluate the zero-shot performance of 30 advanced MLLMs on the 14 subtasks.
+The evaluated MLLMs include BLIP-2 [23], InstructBLIP [12], MiniGPT-4 [59], PandaGPT [39], Multimodal-GPT [15], VisualGLM-6B [5], ImageBind-LLM [17], VPGTrans [53], LaVIN [33], mPLUG-Owl [48], Octopus [3], Muffin [51], Otter [22], LRV-Instruction [28], Cheetor [24], LLaMA-Adapter-v2 [14], GIT2 [41], BLIVA [18], Lynx [52], MMICL [54], GPT-4V [37], Skywork-MM [4], mPLUG-Owl2 [48], Qwen-VL-Chat [9], XComposer-VL [7], LLaVA [29], Lion [2], SPHINX [27], InfMLLM [1], and WeMM [6].
 As displayed in Fig.
@@
 3 also provides a comparison from the other perspective.
-More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination, as shown in Fig.
+We can see the range that current MLLMs can reach in each capability dimension.
+More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination [50, 25], as shown in Fig.
 4.
-These findings are instructive for the subsequent model optimization.
+It is expected that these findings are instructive for the subsequent model optimization.
 In summary, the contributions of this work are as follows: (1) We propose a new benchmark MME to meet the urgent need of MLLM evaluation.
-As far as we know, MME is the first comprehensive MLLM evaluation benchmark [^2] The concurrent works are LAMM-Benchmark Yin et al.
-(2023) and LVLM-eHub Xu et al.
-(2023), which are rather different from our MME.
-We recommend reading the two works as well to enrich the understanding of MLLM evaluation..
-(2) We conduct abundant experiments on a total of 14 subtasks to comprehensively evaluate 10 up-to-date MLLMs.
-(3) We summarize the exposed problems in experiments, proving guidance for the evolution of MLLM.
+(2) A total of 30 up-to-date MLLMs are evaluated on our MME.
+(3) We summarize the exposed problems in experiments, proving guidance for the evolution of MLLMs.
+Model Existence Count Position Color OCR Poseter Celebrity
+ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+
+BLIP-2 86.67 73.33 75.00 60.00 56.67 16.67 81.67 66.67 70.00 40.00 79.25 62.59 68.53 37.06
+mPLUG-Owl 73.33 46.67 50.00 0.00 50.00 0.00 51.67 3.33 55.00 10.00 78.23 57.82 66.18 34.12
+ImageBind-LLM 75.00 53.33 50.00 10.00 43.33 3.33 56.67 16.67 60.00 20.00 52.72 12.24 55.29 21.18
+InstructBLIP 95.00 90.00 80.00 63.33 53.33 13.33 83.33 70.00 57.50 15.00 74.15 49.66 67.06 34.12
+VisualGLM-6B 61.67 23.33 50.00 0.00 48.33 0.00 51.67 3.33 42.50 0.00 53.74 12.24 50.88 2.35
+Multimodal-GPT 48.33 13.33 48.33 6.67 45.00 13.33 55.00 13.33 57.50 25.00 42.86 14.97 49.12 24.71
+PandaGPT 56.67 13.33 50.00 0.00 50.00 0.00 50.00 0.00 50.00 0.00 56.80 19.73 46.47 10.59
+LaVIN 95.00 90.00 61.67 26.67 53.33 10.00 58.33 16.67 67.50 40.00 59.18 20.41 37.94 9.41
+Cheetor 93.33 86.67 63.33 33.33 56.67 23.33 70.00 46.67 65.00 35.00 81.29 65.99 87.65 76.47
+GIT2 96.67 93.33 71.67 46.67 60.00 36.67 85.00 73.33 55.00 10.00 61.56 51.02 81.18 64.71
+GPT-4V 96.67 93.33 86.67 73.33 65.00 30.00 80.00 70.00 95.00 90.00 96.94 95.24 0.00 0.00
+XComposer-VL 96.67 93.33 85.00 73.33 73.33 53.33 88.33 76.67 75.00 50.00 85.71 76.19 83.24 67.06
+LLaVA 95.00 90.00 85.00 70.00 76.67 56.67 90.00 80.00 75.00 50.00 86.39 74.15 83.53 69.41
+LRV-Instruction 88.33 76.67 68.33 43.33 56.67 30.00 88.33 76.67 70.00 40.00 78.77 60.27 67.35 45.29
+Lion 96.67 93.33 85.00 70.00 83.33 70.00 93.33 86.67 57.50 15.00 93.88 87.76 82.94 67.65
+Lynx 98.33 96.67 81.67 70.00 60.00 30.00 90.00 80.00 57.50 20.00 74.49 50.34 71.76 46.47
+MMICL 90.00 80.00 86.67 73.33 55.00 26.67 83.33 73.33 60.00 40.00 81.63 64.63 79.41 62.35
+Muffin 98.33 96.67 86.67 76.67 53.33 13.33 88.33 76.67 52.50 5.00 78.57 59.18 56.47 25.29
+Octopus 93.33 86.67 50.00 3.33 45.00 3.33 66.67 36.67 55.00 10.00 78.23 59.86 75.29 54.12
+Otter 98.33 96.67 58.33 30.00 60.00 26.67 70.00 43.33 57.50 15.00 78.91 59.86 90.88 81.76
+Qwen-VL-Chat 85.00 73.33 83.33 66.67 75.00 53.33 90.00 80.00 80.00 60.00 92.18 86.39 72.35 48.24
+SPHINX 98.33 96.67 86.67 73.33 83.33 70.00 86.67 73.33 62.50 25.00 87.41 76.87 92.65 85.29
+Skywork-$M$ 93.33 86.67 81.67 70.00 46.67 16.67 81.67 63.33 87.50 75.00 91.50 84.35 86.76 73.53
+VPGTrans 56.67 13.33 61.67 23.33 53.33 10.00 56.67 16.67 57.50 20.00 60.88 23.13 51.18 2.35
+WeMM 98.33 96.67 80.00 60.00 73.33 53.33 88.33 80.00 82.50 65.00 86.39 74.15 92.65 86.47
+BLIVA 93.33 86.67 78.33 60.00 58.33 23.33 93.33 86.67 62.50 25.00 84.35 70.75 80.29 60.59
+InfMLLM 96.67 93.33 81.67 70.00 80.00 63.33 95.00 90.00 77.50 55.00 87.07 76.19 86.18 75.29
+LLaMA-AdapterV2 95.00 90.00 73.33 60.00 50.00 6.67 71.67 46.67 67.50 35.00 81.97 65.99 77.94 58.82
+MiniGPT-4 51.67 16.67 48.33 6.67 43.33 0.00 55.00 20.00 52.50 5.00 36.39 5.44 45.00 9.41
+mPLUG-Owl2 95.00 90.00 85.00 70.00 61.67 26.67 83.33 66.67 67.50 35.00 86.73 73.47 87.94 76.47
+Table 1: Evaluation results on the subtasks of existence, count, position, color, OCR, poster, and celebrity. The top two results on each subtask are bolded and underlined, respectively.
+Model Scene Landmark Artwork Comm. Num. Text. Code.
+ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+
+BLIP-2 81.25 64.00 79.00 59.00 76.50 60.00 68.57 41.43 40.00 0.00 55.00 10.00 55.00 20.00
+mPLUG-Owl 78.00 57.50 86.25 73.00 63.25 33.00 57.14 21.43 50.00 10.00 60.00 20.00 47.50 10.00
+ImageBind-LLM 68.75 44.50 53.00 9.00 54.25 16.50 40.00 8.57 50.00 5.00 50.00 0.00 50.00 10.00
+InstructBLIP 84.00 69.00 59.75 20.00 76.75 57.50 75.00 54.29 35.00 5.00 55.00 10.00 47.50 10.00
+VisualGLM-6B 81.75 64.50 59.75 24.00 55.25 20.00 35.00 4.29 45.00 0.00 50.00 0.00 47.50 0.00
+Multimodal-GPT 50.50 17.50 48.25 21.50 46.50 13.00 43.57 5.71 42.50 20.00 50.00 10.00 45.00 10.00
+PandaGPT 72.50 45.50 56.25 13.50 50.25 1.00 56.43 17.14 50.00 0.00 52.50 5.00 47.50 0.00
+LaVIN 78.75 58.00 64.00 29.50 59.25 28.00 58.57 28.57 55.00 10.00 47.50 0.00 50.00 0.00
+Cheetor 84.50 71.50 81.41 64.32 67.50 46.00 64.29 34.29 57.50 20.00 42.50 15.00 57.50 30.00
+GIT2 86.00 72.50 79.50 61.00 81.25 65.00 66.43 32.86 40.00 10.00 52.50 15.00 45.00 0.00
+GPT-4V 83.50 67.50 79.25 59.00 82.00 66.00 79.29 62.86 75.00 55.00 55.00 20.00 90.00 80.00
+XComposer-VL 86.25 73.50 87.75 77.50 73.25 53.00 77.14 61.43 40.00 15.00 67.50 45.00 55.00 30.00
+LLaVA 86.75 74.50 90.00 80.50 70.75 47.00 73.57 54.29 37.50 5.00 57.50 20.00 42.50 5.00
+LRV-Instruction 81.04 65.93 86.84 73.68 63.75 37.50 65.00 35.71 45.00 25.00 55.00 30.00 57.50 15.00
+Lion 85.50 73.50 91.00 82.00 75.75 55.00 74.29 51.43 65.00 40.00 82.50 65.00 42.50 25.00
+Lynx 88.00 76.50 87.00 75.00 72.50 47.00 66.43 44.29 17.50 0.00 42.50 0.00 45.00 0.00
+MMICL 83.75 70.00 76.96 59.16 76.50 59.00 76.43 60.00 47.50 35.00 72.50 60.00 47.50 30.00
+Muffin 83.75 67.50 81.25 65.00 71.00 45.50 68.57 41.43 45.00 0.00 57.50 15.00 47.50 15.00
+Octopus 84.75 72.50 75.00 51.00 64.50 30.50 64.29 35.71 47.50 0.00 62.50 40.00 47.50 15.00
+Otter 86.25 72.50 78.75 58.50 75.00 54.00 66.43 40.00 52.50 20.00 52.50 5.00 55.00 15.00
+Qwen-VL-Chat 83.75 68.50 88.00 76.00 73.50 52.00 76.43 54.29 35.00 5.00 82.50 65.00 32.50 10.00
+SPHINX 86.50 73.50 89.20 78.89 77.00 57.00 75.71 54.29 40.00 15.00 50.00 25.00 50.00 0.00
+Skywork-$M$ 79.29 59.60 75.51 51.53 69.43 45.08 72.14 54.29 65.00 30.00 60.00 20.00 45.00 10.00
+VPGTrans 80.25 61.50 54.75 10.00 57.75 19.50 50.00 14.29 50.00 0.00 57.50 20.00 52.50 5.00
+WeMM 91.75 84.50 90.75 81.50 85.00 71.00 80.00 60.00 47.50 10.00 75.00 55.00 72.50 45.00
+BLIVA 83.50 68.00 63.00 26.50 77.25 56.00 77.86 58.57 47.50 10.00 57.50 20.00 50.00 10.00
+InfMLLM 87.75 77.50 88.50 78.50 68.00 40.50 76.43 55.71 45.00 15.00 67.50 35.00 42.50 10.00
+LLaMA-AdapterV2 85.25 71.00 88.44 79.40 73.25 50.50 67.86 38.57 47.50 0.00 67.50 45.00 60.00 30.00
+MiniGPT-4 54.25 17.50 45.50 8.50 50.00 10.50 46.43 12.86 45.00 0.00 0.00 0.00 40.00 0.00
+mPLUG-Owl2 83.25 70.00 85.75 71.50 77.25 57.00 71.43 44.29 35.00 0.00 67.50 35.00 45.00 15.00
+Table 2: Evaluation results on the subtasks of scene, landmark, artwork, commonsense reasoning, numerical calculation, text translation, and code reasoning.
+The top two results on each subtask are bolded and underlined, respectively.
 2 MME Evaluation Suite
@@
 1.
-When MLLM answers both of these questions correctly, it appears more confident that the MLLM actually comprehends the image and the corresponding knowledge behind it.
+When MLLM answers both of the two questions correctly, it appears more confident that the MLLM actually comprehends the image and the corresponding knowledge behind it, rather than just guessing.
 2.2 Evaluation Metric
@@
 The cognition score is calculated in the same way.
-Figure 2: Leaderboards on our MME benchmark.
-(1) and (2) are the overall leaderboards of perception and cognition respectively, in which the full score of the former is 2000 and that of the latter is 800.
-(3)-(16) are the leaderboards of the 14 subtasks with the full score of 200.
-The score is the sum of the accuracy and the accuracy+ in Tables.
-1, 2, and 3.
-Figure 3: Comparison of 10 advanced MLLMs on 14 subtasks. The full score of each subtask is 200.
+Therefore, the full scores of perception and cognition are 2000 and 800, respectively.
+Figure 3: Comparison of 6 top MLLMs on 14 subtasks. The full score of each subtask is 200.
 2.3 Data Collection
 2.3.1 Perception Tasks
-We argue that perception is one of the most fundamental capabilities of MLLM, and the lack of perception will easily lead to the object hallucination problem Li et al.
-(2023d).
+We argue that perception is one of the most fundamental capabilities of MLLMs, and the lack of perception will easily lead to the object hallucination problem [50, 25].
 That is, MLLM will answer questions based on its own fantasies rather than based on the realistic content of the image, as displayed in Fig.
@@
 The contents of coarse-grained recognition include the existence of common objects, and their count, color, and position.
-The images are sampled from COCO Lin et al.
-(2014), but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations.
+The images are sampled from COCO [26], but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations.
 Even if MLLMs have seen these COCO images, our manually prepared pairs are not presented in their training sets.
-This requires MLLM to be able to understand the instructions and infer corresponding answers.
+This requires MLLMs to be able to understand the instructions and infer corresponding answers.
 In each perception subtask of existence, count, color, and position, we prepare 30 images with 60 instruction-answer pairs.
 Fine-Grained Recognition.
-The fine-grained recognition is more about testing the knowledge resources of MLLM.
+The fine-grained recognition is more about testing the knowledge resources of MLLMs.
 The subtasks consist of recognizing movie posters, celebrities, scenes, landmarks, and artworks, containing 147, 170, 200, 200, and 200 images respectively.
 For the celebrities, we plot a red box to a person with a clearly visible face in the image, and the corresponding instruction is “Is the actor inside the red box named [celebrity name]?
-Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets Huang et al.
-(2020); Mao et al.
-(2017, 2019); Zhou et al.
-(2014); Weyand et al.
-(2020) and all of the instructions are manually designed.
+Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets [19, 34, 35, 58, 44] and all of the instructions are manually designed.
 OCR.
-Optical Character Recognition (OCR) is also a foundational capability of MLLM, serving for subsequent text-based tasks such as text translation and text understanding.
-The images are from Liu et al.
-(2019) and the instruction-answer pairs are manually designed.
-Considering that MLLM is still in its infancy, we only choose the relatively simple samples in this version of MME.
+Optical Character Recognition (OCR) is also a foundational capability of MLLMs, serving for subsequent text-based tasks such as text translation and text understanding.
+The images are sampled from [30] and all of the instruction-answer pairs are manually designed.
+Considering that MLLMs are still in its infancy, we only choose the relatively simple samples in this version of MME.
 The numbers of image and instruction-answer pairs are 20 and 40, respectively.
 2.3.2 Cognition Tasks
-We evaluate if any MLLM can carry out further logical reasoning after perceiving the image, which is the most fascinating aspect of MLLM over previous traditional methods.
-In order to infer the correct answer, MLLM needs to follow the instruction, perceive the contents of the image, and invoke the knowledge reserved in LLM, which is much more challenging than the single perception tasks.
+We evaluate if any MLLM can carry out further logical reasoning after perceiving the image, which is the most fascinating aspect of MLLMs over previous traditional methods.
+In order to infer the correct answer, MLLMs need to follow the instruction, perceive the contents of the image, and invoke the knowledge reserved in LLMs, which is much more challenging than the single perception tasks.
 Examples of the following subtasks are shown in Fig.
@@
 Commonsense Reasoning.
-Unlike the ScienceQA dataset Lu et al.
-(2022) that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life.
-For example, given a photo of a down jacket, asking MLLM whether it is appropriate to wear the cloth when it is cold (or hot).
+Unlike the ScienceQA dataset [32] that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life.
+For example, given a photo of a down jacket, asking MLLMs whether it is appropriate to wear the cloth when it is cold (or hot).
 These are basic knowledge that humans can judge instantly without complex step-by-step reasoning.
-Therefore, we expect MLLM to perform well in a zero-short setting.
+Therefore, we expect MLLMs to perform well in a zero-short setting.
 The images are all manually photographed or generated by diffusion models, and the instruction-answer pairs are all manually designed.
@@
 Numerical Calculation.
-It requires MLLM to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in Huang et al.
-(2023).
-In this version, we only consider relatively easy arithmetic problems, such as addition, subtraction, multiplication, and division.
-There are a total of 20 images and 40 instruction-answer pairs.
+It requires MLLMs to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in [20].
+In this version, we only consider relatively easy arithmetic problems, such as addition and multiplication.
+There are 20 images and 40 instruction-answer pairs.
 The images are all manually taken, and the instruction-answer pairs are all manually designed.
 Text Translation.
-Considering that the MLLM git (2023i) supports both English and Chinese, we set the text translation task.
-It requires MLLM to translate the Chinese written in an image to the corresponding English.
-In this version, we only design basic translation problems, which will be updated according to the development of MLLM in the future.
+Considering that the MLLM [5] supports both English and Chinese, we set the text translation subtask.
+It requires MLLMs to translate the Chinese written in an image to the corresponding English.
+In this version, we only design basic translation problems, which will be updated according to the development of MLLMs in the future.
 The images of this part are all manually taken, and the instruction-answer pairs are all manually designed.
@@
 Code Reasoning.
-It requires MLLM to read the code in the images and automatically complete logical operation inside the code.
-A similar task that writes website code based on an image has been demonstrated in Zhu et al.
-(2023).
+It requires MLLMs to read the code in the images and automatically complete logical operation inside the code.
+A similar task that writes website code based on an image has been demonstrated in [59].
 The images are all manually taken, and the instruction-answer pairs are all manually designed.
@@
 3 Experiments
-Model Existence Count Position Color OCR ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 86.67 73.33 75.00 60.00 56.67 16.67 81.67 66.67 70.00 40.00 LLaVA 50.00 0.00 50.00 0.00 50.00 0.00 51.67 3.33 50.00 0.00 MiniGPT-4 75.00 60.00 66.67 56.67 56.67 33.33 71.67 53.33 62.50 35.00 mPLUG-Owl 73.33 46.67 50.00 0.00 50.00 0.00 51.67 3.33 55.00 10.00 LLaMA-AdapterV2 76.67 56.67 58.33 6.67 43.33 3.33 55.00 16.67 57.50 15.00 InstructBLIP 95.00 90.00 80.00 63.33 53.33 13.33 83.33 70.00 57.50 15.00 VisualGLM-6B 61.67 23.33 50.00 0.00 48.33 0.00 51.67 3.33 42.50 0.00 Otter 53.33 6.67 50.00 0.00 50.00 0.00 51.67 3.33 50.00 0.00 Multimodal-GPT 46.67 10.00 51.67 6.67 45.00 13.33 55.00 13.33 57.50 25.00 PandaGPT 56.67 13.33 50.00 0.00 50.00 0.00 50.00 0.00 50.00 0.00
-Table 1: Evaluation of coarse-grained recognition and OCR.
-Model Poster Celebrity Scene Landmark Artwork ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 79.25 62.59 68.53 37.06 81.25 64.00 79.00 59.00 76.50 60.00 LLaVA 50.00 0.00 48.82 0.00 50.00 0.00 50.00 0.00 49.00 0.00 MiniGPT-4 49.32 19.73 58.82 24.71 68.25 45.50 59.75 30.50 56.25 27.00 mPLUG-Owl 77.89 57.14 66.18 34.12 78.00 57.50 86.25 73.00 63.25 33.00 LLaMA-AdapterV2 52.72 10.88 55.00 21.18 68.75 44.50 53.00 9.00 52.50 14.50 InstructBLIP 74.15 49.66 67.06 34.12 84.00 69.00 59.75 20.00 76.75 57.50 VisualGLM-6B 54.42 12.24 50.88 2.35 81.75 64.50 59.75 24.00 55.75 20.00 Otter 45.24 0.00 50.00 0.00 55.00 14.50 52.00 4.50 48.00 5.50 Multimodal-GPT 45.24 17.01 49.12 24.12 50.50 17.50 50.50 23.00 46.00 12.00 PandaGPT 56.80 19.73 46.47 10.59 72.50 45.50 56.25 13.50 50.25 1.00
-Table 2: Evaluation of fine-grained recognition.
-Model Commonsense Reasoning Numerical Calculation Text Translation Code Reasoning ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ BLIP-2 68.57 41.43 40.00 0.00 55.00 10.00 55.00 20.00 LLaVA 49.29 11.43 50.00 0.00 52.50 5.00 50.00 0.00 MiniGPT-4 58.57 34.29 47.50 20.00 42.50 15.00 67.50 45.00 mPLUG-Owl 59.29 24.29 50.0 10.00 60.00 20.00 47.50 10.00 LLaMA-Adapter V2 54.29 14.29 52.50 5.00 52.50 5.00 52.50 10.00 InstructBLIP 75.00 54.29 35.00 5.00 55.00 10.00 47.50 0.00 VisualGLM-6B 45.71 12.86 45.00 0.00 55.00 10.00 50.00 0.00 Otter 48.57 10.00 47.50 10.00 55.00 10.00 50.00 0.00 Multimodal-GPT 45.71 5.71 50.00 20.00 50.00 5.00 45.00 10.00 PandaGPT 56.43 17.14 50.00 0.00 52.50 5.00 47.50 0.00
-Table 3: Evaluation of cognition.
-In this section, we conduct massive experiments on our MME benchmark to evaluate a total of 10 open-source MLLMs, including BLIP-2, LLaVA, MiniGPT-4, mPLUG-Owl, LLaMA-Adapter-v2, Otter, Multimodal-GPT, InstructBLIP, VisualGLM-6B, and PandaGPT.
-Except BLIP-2, other models have been fine-tuned on their instruction tuning datasets.
-3.1 Models
-All experiments are conducted on NVIDIA V100 GPUs, and we use as large models as possible.
-BLIP-2 Li et al.
-(2023c) focuses on the basic pre-training, where the image encoder and the LLM are all frozen and a light-weight Q-Former is trained for multimodal representation alignment and vision-to-language generation.
-Despite not performing the multimodal instruction tuning, BLIP-2 is still able to follow instructions with the LLM’s own capabilities.
-The version under investigation is “blip2-pretrain-flant5xxl” git (2023a).
-LLaVA Liu et al.
-(2023) is a pioneer in bringing multimodal instruction tuning to MLLM.
-The tuning data are generated by language-only GPT-4 through carefully designed prompts.
-The instructions consist of three types, including conversation, detailed description, and complex reasoning.
-A projection layer and LLM are updated during training.
-The tested version is “LLaVA-7B-v0” git (2023d).
-MiniGPT-4 Zhu et al.
-(2023) performs multimodal instruction tuning on the pre-trained BLIP-2, where only a linear layer is updated.
-The instructions are mainly based on the image caption task, such as “Describe this image in detail.” We use “minigpt4-aligned-with-vicuna13b” git (2023e) for testing.
-mPLUG-Owl Ye et al.
-(2023) uses both language-only instruction data and the multimodal instruction data from LLaVA.
-The visual encoder, the proposed visual abstractor, and LLM (insert a LoRA module Hu et al.
-(2021)) are all updated during training.
-We test “mplug-owl-llama-7b” git (2023f).
-LLaMA-Adapter V2 Gao et al.
-(2023) is only trained on language-only instruction data and image-text pairs via a parameter-efficient manner.
-Additional expert models, such as detection, can be integrated to boost reasoning.
-The version of “LLaMA-7B” git (2023c) is used in testing.
-InstructBLIP Dai et al.
-(2023) is based on BLIP-2 and reorganizes 26 existing public datasets like image caption and VQA as instruction tuning format.
-Only the Q-Former is updated during training.
-“blip2-instruct-flant5xxl” git (2023b) is tested.
-VisualGLM-6B git (2023i) is an open-source MLLM but with unclear training details by now. It supports both English and Chinese. We test “VisualGLM-6B”  git (2023i).
-Otter Li et al.
-(2023b) mixes multimodal in-context learning and multimodal instruction tuning.
-The framework is based OpenFlamingo Awadalla et al.
-(2023), whose perceiver module and parts of LLM are updated during training.
-The instruction data are from VQAv2 Antol et al.
-(2015), GQA Hudson and Manning (2019), LLaVA, and a not public video dataset.
-The version of “OTTER-9B-LA-InContext” git (2023g) is tested.
-Multimodal-GPT Gong et al.
-(2023) is also based on OpenFlamingo and uses both language-only and multimodal instruction data.
-The latter includes data from LLaVA, Mini-GPT4, and various image caption and VQA datasets.
-Only the LoRA module in LLM is updated during training.
-The tested version is “Multimodal-GPT-9B”.
-PandaGPT Su et al.
-(2023) uses multimodal instruction tuning data from LLaVA and MiniGPT-4, and only a linear projection layer as well as the LoRA module in LLM are trainable.
-The pretrained ImageBind Girdhar et al.
-(2023) is adopted as the multimodal encoder, which has an attribute of modality alignment.
-“pandagpt-7b-max-len-512” git (2023h) is used for testing.
+In this section, a total of 30 MLLMs are evaluated on our MME benchmark, including BLIP-2 [23], InstructBLIP [12], MiniGPT-4 [59], PandaGPT [39], Multimodal-GPT [15], VisualGLM-6B [5], ImageBind-LLM [17], VPGTrans [53], LaVIN [33], mPLUG-Owl [48], Octopus [3], Muffin [51], Otter [22], LRV-Instruction [28], Cheetor [24], LLaMA-Adapter-v2 [14], GIT2 [41], BLIVA [18], Lynx [52], MMICL [54], GPT-4V [37], Skywork-MM [4], mPLUG-Owl2 [48], Qwen-VL-Chat [9], XComposer-VL [7], LLaVA [29], Lion [2], SPHINX [27], InfMLLM [1], and WeMM [6].
 3.1 Results
@@
 2 (3)-(6) show the score leaderboards of individual coarse-grained recognition subtasks.
-With respect to the object existence, InstructBLIP gets the highest score 185, with a 95% accuracy and a 90% accuracy+ listed in Table 1.
-Contrastively, the second place BLIP-2 and the third place MiniGPT-4 lag behind InstructBLIP by 25 and 50 scores, respectively.
-For the object count, position, and color, InstructBLIP, BLIP-2, and MiniGPT-4 still make the top three, which is reasonable because the object existence is the basis of object attributes.
-What is interesting is that both InstructBLIP and MiniGPT-4 are instruction finetuned on the pre-trained BLIP-2.
+With respect to the object existence, Otter, Lynx, WeMM, Muffin, and SPHINX get the highest score 195, with a 98.33% accuracy and a 96.67% accuracy+ listed in Table 1.
+Contrastively, the second place, including GIT2, XComposer-VL, Lion, GPT-4V, and etc, lag behind the first place only by 5 scores.
+The results show that these models already have a good performance on object existence.
+For the object count, position, and color, Muffin, Lion (parallel with SPHINX), and InfMLLM make the top one, respectively.
+It suggests that different models have their own strengths.
 Note that in the four coarse-grained subtasks, these MLLMs get the worst results on object position, indicating that the current models are not sensitive enough to the position information.
 Figs.
-2 (7)-(11) show the score leaderboards of individual fine-grained recognition subtasks.
-Regarding to poster recognition, BLIP-2, mPLUG-Owl, and InstructBLIP are the top three.
-It is interesting that mPLUG-Owl underperforms in the coarse-grained recognition, but now it exhibits good.
-This implies that our division of coarse-grained and fine-grained is reasonable and enables us to examine different aspects of MLLMs.
-For the celebrity recognition, BLIP-2, InstructBLIP, and mPLUG-Owl still take the top three with similar scores.
-For the scene recognition, InstructBLIP, VisualGLM-6B, and BLIP-2 ahead of other MLLMs.
-This is the first time VisualGLM-6B has broken into the top three.
-Also noteworthy is PandaGPT, which scores more than a hundred for the first time.
-For the landmark recognition, top three places are taken by mPLUG-Owl, BLIP-2, and MiniGPT-4 respectively, of which mPLUG-Owl gets the top spot.
-For the artwork recognition, BLIP-2, InstructBLIP, and mPLUG-Owl exceed other counterparts, where the first two scores are similar, leading the third by 38+.
+2 (7)-(11) display the score leaderboards of individual fine-grained recognition subtasks.
+Regarding to poster recognition, GPT-4V, Lion, and Qwen-VL-Chat are the top three.
+It is interesting that Qwen-VL-Chat relatively underperforms in the coarse-grained recognition, but now it exhibits good.
+This implies that our division of coarse-grained and fine-grained is reasonable, enabling us to examine different aspects of MLLMs.
+For the celebrity recognition, WeMM, SPHINX, and Otter take the top three with similar scores.
+It is worth noting that GPT-4V refuses to answer questions that involve individuals, resulting in a zero score in the celebrity subtask.
+For the scene recognition, WeMM, InfMLLM, and Lynx ahead of other MLLMs.
+This is the first time InfMLLM and Lynx have broken into the top three in the fine-grained recognition subtasks.
+For the landmark recognition, top three places are taken by Lion, WeMM, and LLaVA respectively, of which Lion gets the top spot.
+For the artwork recognition, WeMM, GPT-4V, and GIT2 exceed other counterparts, where the last two scores are similar.
+Note that GPT-4V declines to answer some questions about private art collection, which lowers its score.
 With respect to OCR listed in Fig.
-2 (12), BLIP-2, MiniGPT-4, and Multimodal-GPT get the top three with scores of 110, 97.5, and 82.5 respectively, in which the first two perform more outstanding.
-Figure 4: Common problems revealed in our experiments. [Y]/[N] means the ground truth answer is yes/no. [R] is the generated answer of MLLM.
-As can be seen in Fig.
-2 (1), in the leaderboard of the whole perception recognition, BLIP-2, InstructBLIP, and MiniGPT-4 come in top three, closely followed by mPLUG-Owl.
-Their perception scores are 1293.84, 1212.82, 1010.66, and 966.32, respectively.
-Among them, mPLUG-Owl shows a clear bias, good at fine-grained recognition, but not good at coarse-grained recognition.
-On the contrary, MiniGPT-4 has worse fine-grained recognition than coarse-grained recognition, while the discrepancy is not that significant as mPLUG-Owl.
-Comparatively, the overall performances of BLIP-2 and InstructBLIP are more stable.
+2 (12), GPT-4V, Skywork-MM, and WeMM get the top three with scores of 185, 162.5, and 147.5 respectively.
+GPT-4V presents a huge advantage, leading the other two models by 22+ socres.
+As presented in Fig.
+2 (1), in the leaderboard of the whole perception recognition, WeMM, InfMLLM, and SPHINX come in top three, closely followed by Lion, LLaVA, and XComposer-VL.
+Figure 4: Common problems revealed in experiments. [Y]/[N] means the ground truth answer is yes/no. [R] is the generated answer.
 3.1.2 Cognition
@@
 2 (13)-(16) plot the score leaderboards of individual subtasks.
-In terms of the commonsense reasoning, the “ever-victorious generals” InstructBLIP, BLIP-2, and MiniGPT-4 exceed other MLLMs, especially InstructBLIP, which gets a score of 129.29.
-With respect to numerical calculation and text translation, although the designed questions are moderate in difficulty, these MLLMs do not perform satisfactorily, with no score beyond 80.
+In terms of the commonsense reasoning, the “ever-victorious generals” GPT-4V, WeMM, and XComposer-VL exceed other MLLMs, especially GPT-4V, which gets a score of 142.14.
+With respect to numerical calculation, GPT-4V still achieves first place, but falls short in the text translation.
+Regardless of whether it is commonsense reasoning, numerical calculation, or text translation, none of the highest scores exceed 150.
 This suggests that MLLMs have a lot of room for improvement in these capabilities.
-For the code reasoning, MiniGPT-4 achieves a high score of 112.5, far ahead of other counterparts.
-This is reminiscent of the example it showed of writing website codes from an image.
-For all of the cognition tasks, MiniGPT-4, InstructBLIP, and BLIP-2 win the gold, silver, and bronze medals respectively, as shown in Fig.
+For the code reasoning, GPT-4V achieves a high score of 170, far ahead of other counterparts.
+For all of the cognition tasks, GPT-4V, Lion, and WeMM win the gold, silver, and bronze medals respectively, as shown in Fig.
 2 (2).
@@
 These phenomena indicate that the logic chain is broken during the reasoning process of MLLMs.
-Adding CoT prompts, such as “Let’s think step by step” Driess et al.
-(2023), may yield better results.
+Adding CoT prompts, such as “Let’s think step by step” [13], may yield better results.
 We look forward to a further in-depth research.
@@
 When the instruction contains descriptions of an object that does not appear in the image, the MLLM will imagine that the object exists and ultimately gives a “yes” answer.
-Such a case of constantly answering “yes” results in an accuracy about 50% and an accuracy+ about 0, as shown in Tables 1, 2, and 3.
+Such a case of constantly answering “yes” results in an accuracy about 50% and an accuracy+ about 0, as shown in Tables 1 and 2.
 This suggests an urgent need to suppress hallucinations, and the community should take into account of the reliability of the generated answers.
@@
 This paper has presented the first MLLM evaluation benchmark MME that has four distinct characteristics in terms of task type, data source, instruction design, quantitative statistics.
-We evaluate 10 advanced MLLMs on MME and the experimental results show that there is still a large room to improve.
+30 advanced MLLMs are evaluated on MME and the experimental results show that there is still a large room to improve.
 We also summarize the common problem raised in experimental results, providing valuable guidance for the development of MLLM.

### v5

@@
 However, it is difficult for these case studies to fully reflect the performance of MLLM, lacking a comprehensive evaluation.
-In this paper, we fill in this blank, presenting the first comprehensive MLLM Evaluation benchmark MME [^1] MME is collected by Xiamen University for academic research only.
-This is the v1 version and will be updated as MLLMs evolve..
+In this paper, we fill in this blank, presenting the first comprehensive MLLM Evaluation benchmark MME.
 It measures both perception and cognition abilities on a total of 14 subtasks.
@@
 A total of 30 advanced MLLMs are comprehensively evaluated on our MME, which not only suggests that existing MLLMs still have a large room for improvement, but also reveals the potential directions for the subsequent model optimization.
-The data application manner and online leaderboards are released at https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation.
+The data are released at the project page: https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation.
 1 Introduction
-The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) [37, 8, 20, 23, 13, 9].
-It refers to using LLM as a brain to process multimodal information and give reasoning results [49].
-Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks [13, 45, 38, 27, 54, 31].
-The three representative abilities of LLM [55], including instruction following [40], In-Context Learning (ICL) [10], and Chain-of-Thought (CoT) [43] are also manifested in multimodality.
-For example, Flamingo [8] turns on multimodal ICL, which can adapt to new tasks by giving a few examples.
-PaLM-E [13] achieves amazing OCR-free math reasoning via CoT.
-GPT-4V [37] shows even more ability in a variety of complex reasoning tasks [46].
-MiniGPT-4 [59] implements GPT-4[37]-like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning.
+The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) openai2023gpt4 ; huang2023language ; li2023blip ; driess2023palm ; team2025kimi ; tu2024overview .
+It refers to using LLM as a brain to process multimodal information and give reasoning results yin2023survey .
+Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks driess2023palm ; wen2025dycrowd ; shen2023hugginggpt ; lin2023sphinx ; zhang2025speechact ; zhao2023mmicl ; liu2023mmbench ; liu2021geometry ; zhang2025logavatar ; ying2024mmt ; fang2024mmbench .The three representative abilities of LLM zhao2023survey , including instruction following touvron2023llama , In-Context Learning (ICL) brown2020language , and Chain-of-Thought (CoT) wei2022chain are also manifested in multimodality.
+For example, Flamingo alayrac2022flamingo turns on multimodal ICL, which can adapt to new tasks by giving a few examples.
+PaLM-E driess2023palm achieves amazing OCR-free math reasoning via CoT.
+GPT-4V openai2023gpt4 shows even more ability in a variety of complex reasoning tasks wu2023early .
+MiniGPT-4 zhu2023minigpt implements GPT-4openai2023gpt4 -like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning.
 These emergent abilities of MLLMs are exciting and imply that a new dawn has broken in artificial intelligence.
@@
 (3)-(16) are the leaderboards of the 14 subtasks with the full score of 200.
-The score is the sum of the accuracy and the accuracy+ in Tables.
-1 and 2.
+The score is the sum of the accuracy and the accuracy+ in Tables 1 and 2.
 A total of 30 advanced MLLMs joint the leaderboards.
@@
 The existing three common quantitative evaluation manners for MLLMs have their limitations that are difficult to comprehensively evaluate performance.
-Specifically, the first manner [47, 12, 42] evaluates on existing traditional multimodal datasets, such as image caption [11] and VQA [16, 36, 32].
+Specifically, the first manner xu2022multiinstruct ; dai2023instructblip ; wang2023visionllm evaluates on existing traditional multimodal datasets, such as image caption chen2015microsoft and VQA goyal2017making ; marino2019ok ; lu2022learn .
 However, on the one hand, it may be hard to reflect the emergent abilities of MLLMs on these datasets.
 On the other hand, since the training sets of large models are no longer unified, it is difficult to guarantee that all MLLMs have not used the testing set for training.
-The second manner [48] is to collect data for an open-ended evaluation, but either the data is unavailable to public by now [57] or the amount is small (only 50 images) [48].
-The third manner focuses on one aspect of MLLMs, such as object hallucination [25] or adversarial robustness [56], which is powerless to comprehensive evaluation.
+The second manner ye2023mplug is to collect data for an open-ended evaluation, but either the data is unavailable to public by now zhao2023chatbridge or the amount is small (only 50 images) ye2023mplug .
+The third manner focuses on one aspect of MLLMs, such as object hallucination li2023evaluating or adversarial robustness zhao2023evaluating , which is powerless to comprehensive evaluation.
 In light of these concerns, a new comprehensive evaluation benchmark is urgently needed to match the flourish of MLLMs.
@@
 The open-ended answer of MLLMs poses significant challenges to the quantization.
-Existing methods tend to use GPT or manual scoring [21, 29, 48], but there may be problems of inaccuracy and subjectivity.
+Existing methods tend to use GPT or manual scoring li2023mimic ; liu2023visual ; ye2023mplug , but there may be problems of inaccuracy and subjectivity.
 To this end, we collect a comprehensive MLLM Evaluation benchmark, named as MME, which meets the above four characteristics at the same time:
@@
 We conduct massive experiments to evaluate the zero-shot performance of 30 advanced MLLMs on the 14 subtasks.
-The evaluated MLLMs include BLIP-2 [23], InstructBLIP [12], MiniGPT-4 [59], PandaGPT [39], Multimodal-GPT [15], VisualGLM-6B [5], ImageBind-LLM [17], VPGTrans [53], LaVIN [33], mPLUG-Owl [48], Octopus [3], Muffin [51], Otter [22], LRV-Instruction [28], Cheetor [24], LLaMA-Adapter-v2 [14], GIT2 [41], BLIVA [18], Lynx [52], MMICL [54], GPT-4V [37], Skywork-MM [4], mPLUG-Owl2 [48], Qwen-VL-Chat [9], XComposer-VL [7], LLaVA [29], Lion [2], SPHINX [27], InfMLLM [1], and WeMM [6].
+The evaluated MLLMs include BLIP-2 li2023blip , InstructBLIP dai2023instructblip , MiniGPT-4 zhu2023minigpt , PandaGPT su2023pandagpt , Multimodal-GPT gong2023multimodal , VisualGLM-6B git-visualglm-6b , ImageBind-LLM han2023imagebind , VPGTrans zhang2023transfer , LaVIN luo2023cheap , mPLUG-Owl ye2023mplug , Octopus git-octopus , Muffin yu2023reformulating , Otter li2023otter , LRV-Instruction liu2023aligning , Cheetor li2023finetuning , LLaMA-Adapter-v2 gao2023llama , GIT2 wang2022git , BLIVA hu2023bliva , Lynx zeng2023matters , MMICL zhao2023mmicl , GPT-4V openai2023gpt4 , Skywork-MM git-skywork , mPLUG-Owl2 ye2023mplug , Qwen-VL-Chat bai2023qwen , XComposer-VL git-xcomposer , LLaVA liu2023visual , Lion git-lion , SPHINX lin2023sphinx , InfMLLM git-infmllm , and WeMM git-wemm .
 As displayed in Fig.
@@
 We can see the range that current MLLMs can reach in each capability dimension.
-More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination [50, 25], as shown in Fig.
+More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination yin2023woodpecker ; li2023evaluating , as shown in Fig.
 4.
@@
 2.3.1 Perception Tasks
-We argue that perception is one of the most fundamental capabilities of MLLMs, and the lack of perception will easily lead to the object hallucination problem [50, 25].
+We argue that perception is one of the most fundamental capabilities of MLLMs, and the lack of perception will easily lead to the object hallucination problem yin2023woodpecker ; li2023evaluating .
 That is, MLLM will answer questions based on its own fantasies rather than based on the realistic content of the image, as displayed in Fig.
@@
 The contents of coarse-grained recognition include the existence of common objects, and their count, color, and position.
-The images are sampled from COCO [26], but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations.
+The images are sampled from COCO lin2014microsoft , but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations.
 Even if MLLMs have seen these COCO images, our manually prepared pairs are not presented in their training sets.
@@
 For the celebrities, we plot a red box to a person with a clearly visible face in the image, and the corresponding instruction is “Is the actor inside the red box named [celebrity name]?
-Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets [19, 34, 35, 58, 44] and all of the instructions are manually designed.
+Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets huang2020movienet ; mao2017deepart ; mao2019visual ; zhou2014learning ; weyand2020google and all of the instructions are manually designed.
 OCR.
 Optical Character Recognition (OCR) is also a foundational capability of MLLMs, serving for subsequent text-based tasks such as text translation and text understanding.
-The images are sampled from [30] and all of the instruction-answer pairs are manually designed.
+The images are sampled from liu2019curved and all of the instruction-answer pairs are manually designed.
 Considering that MLLMs are still in its infancy, we only choose the relatively simple samples in this version of MME.
@@
 Commonsense Reasoning.
-Unlike the ScienceQA dataset [32] that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life.
+Unlike the ScienceQA dataset lu2022learn that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life.
 For example, given a photo of a down jacket, asking MLLMs whether it is appropriate to wear the cloth when it is cold (or hot).
@@
 Numerical Calculation.
-It requires MLLMs to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in [20].
+It requires MLLMs to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in huang2023language .
 In this version, we only consider relatively easy arithmetic problems, such as addition and multiplication.
@@
 Text Translation.
-Considering that the MLLM [5] supports both English and Chinese, we set the text translation subtask.
+Considering that the MLLM git-visualglm-6b supports both English and Chinese, we set the text translation subtask.
 It requires MLLMs to translate the Chinese written in an image to the corresponding English.
@@
 It requires MLLMs to read the code in the images and automatically complete logical operation inside the code.
-A similar task that writes website code based on an image has been demonstrated in [59].
+A similar task that writes website code based on an image has been demonstrated in zhu2023minigpt .
 The images are all manually taken, and the instruction-answer pairs are all manually designed.
@@
 3 Experiments
-In this section, a total of 30 MLLMs are evaluated on our MME benchmark, including BLIP-2 [23], InstructBLIP [12], MiniGPT-4 [59], PandaGPT [39], Multimodal-GPT [15], VisualGLM-6B [5], ImageBind-LLM [17], VPGTrans [53], LaVIN [33], mPLUG-Owl [48], Octopus [3], Muffin [51], Otter [22], LRV-Instruction [28], Cheetor [24], LLaMA-Adapter-v2 [14], GIT2 [41], BLIVA [18], Lynx [52], MMICL [54], GPT-4V [37], Skywork-MM [4], mPLUG-Owl2 [48], Qwen-VL-Chat [9], XComposer-VL [7], LLaVA [29], Lion [2], SPHINX [27], InfMLLM [1], and WeMM [6].
+In this section, a total of 30 MLLMs are evaluated on our MME benchmark, including BLIP-2 li2023blip , InstructBLIP dai2023instructblip , MiniGPT-4 zhu2023minigpt , PandaGPT su2023pandagpt , Multimodal-GPT gong2023multimodal , VisualGLM-6B git-visualglm-6b , ImageBind-LLM han2023imagebind , VPGTrans zhang2023transfer , LaVIN luo2023cheap , mPLUG-Owl ye2023mplug , Octopus git-octopus , Muffin yu2023reformulating , Otter li2023otter , LRV-Instruction liu2023aligning , Cheetor li2023finetuning , LLaMA-Adapter-v2 gao2023llama , GIT2 wang2022git , BLIVA hu2023bliva , Lynx zeng2023matters , MMICL zhao2023mmicl , GPT-4V openai2023gpt4 , Skywork-MM git-skywork , mPLUG-Owl2 ye2023mplug , Qwen-VL-Chat bai2023qwen , XComposer-VL git-xcomposer , LLaVA liu2023visual , Lion git-lion , SPHINX lin2023sphinx , InfMLLM git-infmllm , and WeMM git-wemm .
 3.1 Results
@@
 These phenomena indicate that the logic chain is broken during the reasoning process of MLLMs.
-Adding CoT prompts, such as “Let’s think step by step” [13], may yield better results.
+Adding CoT prompts, such as “Let’s think step by step” driess2023palm , may yield better results.
 We look forward to a further in-depth research.
@@
 30 advanced MLLMs are evaluated on MME and the experimental results show that there is still a large room to improve.
-We also summarize the common problem raised in experimental results, providing valuable guidance for the development of MLLM.
+We also summarize the common problem raised in experimental results, providing valuable guidance.
+Although MME is a comprehensive benchmark, it still has room for improvement in terms of capability coverage, such as covering more scenarios that require reasoning.
+We will continue to iterate MME series in the future to meet the evaluation requirements of MLLMs.
+More importantly, we hope that through the design of benchmarks, we can reflect the thoughts on the next capabilities of the model and thereby promote its development.
+Acknowledgments
+This work is funded by National Natural Science Foundation of China (Grant No.
+62506158 and No.
+62441234), Fundamental Research Funds for the Central Universities, AI & AI for Science Project of Nanjing University (No.
+2024300529), and CCF-Tencent Rhino-Bird Open Research Fund.